Gail works in a flower shop, where she produces 10 floral arrangements per hour. She is paid $10 per hour for the first eight hours she works and $15 an hour for each additional hour she works. What is the firm’s cost function? What are its AC, AVC, and MC functions? Draw the AC, AVC, and MC curves.



1
Expert's answer
2022-02-13T11:51:34-0500

Cost function "C(x)"

"C(x):Total\ Cost= FC+VC"

"FC=\$10\times 8=\$80"

"VC=\%15 \times h=\$15h"

"\therefore C(x)= 80+15h"

Average Cost

"AC=\frac{TC}{h}"

Q here is equivalent to 10 floral arrangements produced within an hour.

"\therefore AC=(80+15h)/h"

"AC=(\frac{80}{h}+15)"

Average Variable Cost

"AVC= \frac{VC}{h} = \frac{15h}{h}=15"

Marginal Cost

to get the MC, we find the partial derivative of the Total cost thereby maintaining 80 as a constant.

"MC= \frac{\delta TC}{\delta h}=80+15=95"

Graphing the functions we have;